[Brahms-l] Shift report 20040115 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Thu Jan 15 2004 - 07:59:44 EST
------------------------------------------------------------
Date: Thu Jan 15 07:59:43 2004, Shift: 00:00-08:00
Supervisor: Roy WADA, with: Ionut ARSENE
Log book#15 pages 33 - 33

Shift Summary:
------------------------------------------------------------
1. Beam from the beggining of the shift until 02:45, and the from 06:20 until the end of the shift.

2. Data runs:
   9678 - MRS@90deg D5@A2255 and FS@4deg D1@B1723 - 200k trig6
   9679 - MRS@90deg D5@A2255 and FS@4deg D1@B1219 - 170k
   9680 - same settings - 50k
   9687 - same settings - 170k
   9688 - same settings - 160k
   Pedestals
   9681 - Non TPC
   9682 TPC
3. MRS trigger was off for all runs and TPM1 is recovered.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jan 15 07:59:54 2004

This archive was generated by hypermail 2.1.8 : Thu Jan 15 2004 - 08:00:12 EST